OpenStudy (anonymous):

Use synthetic division to determine the quotient and remainder (x^3 – 2x^2 – 8x + 5) ÷ (x – 1)

OpenStudy (anonymous):

\[x^2-x-9-\frac{4}{x-1} \]

OpenStudy (dumbcow):

x-c = x-1 --> c=1 1 | 1 -2 -8 5 1 -1 -9 -------------- 1 -1 -9 -4 -4 is the remainder. quotient = x^2 -x -9